1/3x-5/6y=-6
4x+7y=30
What are x and y
Thank you

Answer by mananth(16949)

1/3 x-5/6 y=-6
(2y-5x)/6=-6
2y-5x=-36------------1
4x+7y=30.............2
.....
-5x+2y=-36.............1
4x+7y=30 .............2

multiply (1)by -7
Multiply (2) by 2
35x-14y= 252
8x+14y=60

Add the two equations
43x+y =312
/43
x=7 1/4

plug value of x in (1)
-5x+2y=-36
-36.28+2 y=-36
2y=-36+36.28
2y=0.28
y= 1/7